### v1

@@ 3.3 Pre-training Objective
In ecological contexts, temporal difference learning can be beneficial. Instead of predicting $x_{t+1}$ directly, we predict the increment $\Delta x_{t}=x_{t+1}-x_{t}$. This approach, often encountered in reinforcement learning [31], can reduce biases from unobserved global offsets or stable large-scale means. For instance, daily vegetation changes or seasonal fluctuations in species population can be smaller and more stable than absolute population numbers. By focusing on differences, we encourage the model to learn transition dynamics or, more specifically, biodiversity dynamics:
$\mathcal{L}_{TD}=||\hat{\Delta x_{t}^{v}}-(x_{t+1}^{v}-x_{t}^{v})||$ (6)
This choice is reinforced by empirical results in specific biodiversity modelling tasks (e.g., ephemeral wetlands or short-lived insect populations), which show improved forecasting stability over standard next-step MSE [32].
Going one step forward, given a system state $\mathbf{X}_{t}$ at time $t$, we aim to predict the next state $\mathbf{X}_{t^{\prime}}$ at time $t^{\prime}>t$. In the common single-step forecast scenario, following [18], we define a simulator function
@@ 3.4 Fine-tuning
BioAnalyst is fine-tuned in three different settings, each contributing to a distinct goal. In this section, a description of each setting is provided, and in the next section, we present the quantitative results.
3.4.1 Short-lead time finetuning
@@
In this setting, we follow a similar approach to [18] and fine-tune the entire BioAnalyst for six rollout steps, effectively predicting biodiversity dynamics six months ahead. We freeze the whole architecture, including the encoder, decoder, and backbone, while training only the newly added VeRA adapters [33] on the backbone’s attention heads. We found VeRA to perform equally or sometimes slightly better than other Parameter Efficient Fine-tuning Techniques (PEFTs), such as LoRA [34], which use only one-tenth of the learnable parameters.
3.4.2 Roll-out finetuning
In this setting, following the same technique as before, we increase the trajectory length to 12 monthly observations, effectively predicting biodiversity dynamics one year in advance. The multi-step objective below is used for both short-lead and rollout fine-tuning settings.
$\mathcal{L}_{FT}=\frac{1}{K}\sum_{k=1}^{K}\mathcal{L}_{TD}(\hat{x}^{u}_{t+K},x^{u}_{t+K})$ (9)
The model is rolled forward $K$ time steps, where $K<horizon$, the loss is averaged across steps, and only the last step’s gradients are propagated due to the size of the model, following the push-forward trick [35].
3.4.3 Task specific finetuning
To evaluate the ecological capacity and environmental structure encoded in the BioAnalyst Foundation Model, we implement two complementary fine-tuning tasks. These tasks are designed to interrogate distinct dimensions of the model’s representation space: its ability to adapt to biotic presence-only data under temporal shift, and its capacity to retain structured abiotic gradients related to seasonal climate. For comparison, we also ran these fine-tuning tasks using the Aurora-0.25 model [18], which shares the same model architecture but has been pre-trained on climate modalities only.
@@ 3.4.3 Task specific finetuning
The first task involves partial model adaptation: the BioAnalyst’s encoder and decoder are frozen. At the same time, the backbone is fine-tuned with VeRA adapters using historical species plant occurrence data from the GeoLifeCLEF2024 benchmark dataset [36] to forecast time-series distributions. The second task is diagnostic: a regression head is trained on top of frozen decoder embeddings for both BioAnalyst and Aurora, to predict monthly climate variables from CHELSA v2.1 [37]. Together, these tasks provide a dual lens on the model’s ecological generalisation and environmental coherence.
3.4.4 Biotic fine-tuning: forecasting species distributions
The model takes as input a species distribution matrix at time $t$ and is trained to predict the distribution at time $t+1$. The input and target distributions are normalised using species-specific statistics (mean and standard deviation). The model preserves the full spatial resolution of the data, working with grid-based representations rather than individual occurrence points. Training is performed end-to-end using a combination of loss functions, including the GeoLifeClef defined F1 score [38] and root mean square error (RMSE). The custom F1 score is determined in Equation 17. These metrics indicate whether the model learns both local and global distribution patterns, and also gives a comparable score for the GeoLifeCLEF benchmark.
BioAnalyst’s latent representations are then analysed through Principal Component Analysis (PCA) computation to understand how it encodes species-specific patterns. Visualising PCA helps assess whether the model learns meaningful environmental-species associations that can transfer across time points. This task simulates a realistic ecological forecasting scenario under observational uncertainty and evaluates whether the model’s environmental representation can be adapted to capture transferable species–environment associations.
3.4.5 Abiotic linear probing: recovering seasonal climate structure
Figure 3: Comparison between CHELSA target climate fields and decoded outputs from the BioAnalyst Small Decoder for May 2018 over Europe. Panels (a) and (b) show the downsampled CHELSA reference for precipitation (kg $m^{2}$) and temperature (K/10), respectively. Panels (c) and (d) show the corresponding decoded predictions from the model after reconstruction. The decoded precipitation captures major spatial gradients and orographic patterns (e.g., Alps, Norway), though with some smoothing. The decoded temperature field accurately reproduces latitudinal and coastal gradients present in the target. All maps are normalised for comparison.
@@ 4.1 Rollouts: Forecasting biodiversity dynamics
In this part, we present four main results produced from our pre-training and rollout fine-tuning experiments. First, we present the cumulative mean for all the species distributions, comparing the ground truth distributions with the predicted distributions, considering only their absolute numerical values on Figure 4. The predictions closely follow the data trend, capturing both upward and downward changes in the species distributions which validates BioAnalyst’s capacity to effectively capture species distribution patterns.
Second, we report strong Sørensen similarity score of 0.31 for 28 species across all land grid cells which means that the predicted assemblage reproduces roughly one-third of the observed community patterns. Figure 5 highlights various patches that coincide with densely sampled or under-sampled regions. Overall, the spatial pattern confirms moderate compositional skill, useful for broad biogeographic inference, yet leaving room for improving species co-occurrence dynamics. Third, Figure 6 depicts BioAnalyst ability to localise and spatially predict granular species distributions.
Finally, we highlight the temporal drift of BioAnalyst’s species distribution predictions when performing 12 rollout steps on Figure 11. The species 1898286, 2491534, 8077224 and 9809229 exhibit the highest MAE while all the rest exhibit variable behaviour with the majority increasing the MAE on every next step.
Figure 4: The cumulative mean of all the species distributions in a 12 step rollout, predicting 1 year ahead.
Figure 5: The community Sørensen similarity map produced using Equation 19 with a mean value of 0.31 for the species variable group. Warm shades on the map highlight hotspots where more than half the species are matched, while cool blues mark cells with little or no overlap.
a) Refer to caption
b) Refer to caption
Figure 6: (a) Ground truth and prediction spatial plots for the species Pieris brassicae (ID 1920506) on 01-4-2019 and MAE of 0.00003. (b) Zoomed in plot, highlighting areas of interest, where the model can capture the general distribution, although failing to capture high-density areas.
@@ 4.2 Biotic fine-tuning: forecasting species distributions
To assess the biotic predictive capacity of BioAnalyst’s pre-trained embeddings, we fine-tuned a classification head to forecast species presence in the GeoLifeCLEF 2021 dataset. The model achieved high accuracy with an F1 score of 0.9964 and an RMSE of 0.5284, demonstrating its ability to support fine-scale species distribution modelling. In the same setting, Aurora performs marginally worse in evaluation metrics than BioAnalyst, as shown in Table 1. Nevertheless, investigating the predicted species richness in more detail for both models, we found that Aurora overestimates its predictions, while BioAnalyst is more spatially accurate, as shown in Figure 7. This is an expected result, as BioAnalyst has been trained on 28 species distributions and exhibits better spatial grounding.
Figure 7: Comparison of observed and predicted species richness in the Netherlands for the year 2021. a) The Observed Richness based on empirical field data from the GeoLifeCLEF2024 survey. b) The species richness for 2021 predicted by the BioAnalyst, and c) the predictions from the Aurora Model for 2021. The results have been masked for land only, as both models predicted richness for most ocean pixels as well.
We further analysed the learned representation structure via Principal Component Analysis (PCA) on the backbone outputs. The first three principal components accounted for 96.4% of the total variance, with PC1 alone explaining 54.44%, followed by PC2 (31.05%) and PC3 (10.92%) (Figure 8). This concentration of variance in a few dimensions suggests that the BioAnalyst latent space captures most biotic information in a compact and low-dimensional manifold, facilitating robust downstream generalisation across ecological prediction tasks.
Figure 8: Correlation matrix of the first six principal components (PCs) derived from the BioAnalyst backbone in the GeoLifeCLEF task. The near-zero off-diagonal values indicate that the principal components are orthogonal, confirming effective dimensional separation in the learned latent space.
Table 1: Performance of the species distribution forecasting task predicting species distribution for 2021.
Model Loss F1 RMSE
BioAnalyst 0.0057 0.9964 0.5284
Aurora 0.0130 0.9945 0.5014
@@ 4.3 Abiotic linear probing: recovering seasonal climate structure
The model achieved strong predictive performance, with the best epoch reaching an $R^{2}$ of 0.9002, a loss of 0.0225, and an RMSE of 0.1499, as shown in Table 2. These values indicate that the BioAnalyst decoder outputs contain sufficient information to reconstruct fine-grained seasonal climate patterns even without any fine-tuning of the encoder. In the same theme, comparing BioAnalyst with the pre-trained Aurora-025 yielded a stronger predictive capacity for BioAnalyst. However, this comparison is not entirely representative, as it provides a performance baseline, since BioAnalyst is trained at a monthly frequency, while Aurora is trained at a 6-hour interval.
To further examine the structure of the learned embeddings, we applied Principal Component Analysis (PCA) to the decoder outputs. The resulting correlation matrix of the first six principal components revealed strong orthogonality between components, as expected Figure 9. This suggests that the BioAnalyst representations organise abiotic variation along separable axes of climate seasonality and geography, supporting their use in downstream ecological tasks.
Figure 9: Monthly aggregated correlation matrix of the first six principal components (PCs) derived from the BioAnalyst decoder outputs in the CHELSA linear probing task from 2000-2019. The near-zero off-diagonal values indicate that the principal components are orthogonal, confirming effective dimensional separation in the learned latent space. This supports the interoperability of downstream regression analyses along independent axes of climate variability.
Table 2: Performance of the linear probing task predicting down-sampled CHELSA v2.1 temperature and precipitation targets from BioAnalyst decoder outputs over Europe.
Model Loss R2 RMSE
BioAnalyst 0.0225 0.9002 0.1499
Aurora 0.2668 0.7354 0.5144
@@ 5 Conclusion & Discussion
In this work, we introduced BioAnalyst, the first Foundation Model for biodiversity. BioAnalyst is truly multi-modal, light-weight and can be used to model various complex ecological phenomena with competitive performance, setting a new accuracy baseline for ecological forecasting. We highlighted its predictive analytics as a stand-alone model in tasks such as biodiversity and dynamics modelling, as well as in tasks like species distribution forecasting, absence detection, and monthly climate linear probing.
The results demonstrate that Foundation Models pre-trained on heterogeneous ecological and environmental data can generalise well across a range of predictive biodiversity tasks. Notably, BioAnalyst scales to regional or continental domains without incurring the computational costs typically associated with larger geospatial models. The model’s strong performance in both biotic and abiotic tasks suggests that it learns a unified latent representation of environmental structure that is transferable and biologically meaningful. This opens new opportunities for rapid model adaptation in data-poor contexts and for advancing hypothesis-driven ecological modelling through representation learning. In addition, we have utilised open, available, and licensed data. We have also open-sourced the model weights, training routine, scaling recipe and fine-tuning techniques used during our experiments. We believe that the research community can be greatly benefited by using our pre-trained weights but also inspired by our open-sourced solution to pursue new avenues of research in complex fields with multi-modal data, complex modelling and scalability demands.
The avenues for improvement are numerous, as BioAnalyst has set the proving ground for the application of Foundation Models in ecology, specifically in biodiversity. Future extensions may include incorporating additional modalities, such as genomic or functional trait data, to enable even more expressive and interpretable biodiversity analytics. At present, BioAnalyst produces exclusively deterministic forecasts. Adopting a probabilistic framework is very important for variables that behave stochastically, like species geographic distributions, surface latent-heat flux, precipitation, convective snowfall rate, 10m wind, and more [39]. Towards that end, a promising line of future work could investigate ensembles of BioAnalyst models trained on different datasets or fine-tuning the model into a probabilistic variant.
Another promising avenue for improvement would be the generation of biodiversity-related modalities that stem from EO FMs, such as TerraMind, and continue to fine-tune them for new tasks. It would be interesting to see the correlations between these modalities and the raw modalities used in the pre-training of BioAnalyst. Additionally, expanding BioAnalyst’s capabilities with user interaction features, such as chat, would enhance its interpretability.
However, some caution is warranted when interpreting downstream predictions. For example, while BioAnalyst correctly captures trends such as species distribution decline in the data (Figure 4), this may partly reflect temporal biases in observation effort rather than actual ecological change. Such artefacts highlight the need for careful disentanglement of signal and sampling in biodiversity datasets when training and evaluating foundation models. Future work could address this by incorporating observation effort metadata or by explicitly modelling detection processes alongside ecological predictors.
@@
A key open question for any AI-based modelling is the quantification of uncertainty in the model’s predictions. Adding extra dimensions, such as spatial and temporal, makes the task even more challenging. Methods for quantifying this uncertainty at different granularity levels, such as cartograms [40] or meta-model traits [41], are part of BioAnalyst’s future work. Truly synergistic models that embed ecological principles (e.g. energy budgets, trophic interactions) into AI architectures that simulate population dynamics are an exciting frontier [42]. This could mean neural networks that respect mass-balance constraints or reinforcement learning agents that simulate animal foraging behaviour. Such integration would yield models that not only predict well but also adhere to known ecological laws, making them more generalisable, trustworthy, and aligned with global biodiversity goals [43].
Like any other advanced AI model, BioAnalyst has limitations. One of the biggest is its constrained area of operation, Europe, which is not representative of global biodiversity dynamics. Additionally, BioAnalyst is trained on data that represent biodiversity only at the terrestrial level. It does not take into consideration the sea, for example, amplifying the bias towards other parts of biodiversity and the compound effects they may have on both global and local biodiversity dynamics. Higher-quality, more frequently sampled, and better-curated data points could further enhance BioAnalyst’s capabilities in combination with longer training. As mentioned earlier, the uncertainty quantification of BioAnalyst’s prediction is a project in itself and was not the focus of this work.
Declarations
[…]
Supplementary information
A supplementary file accompanies this article.
[…]
Acknowledgements

### v2

@@ 3.3 Pre-training Objective
In ecological contexts, temporal difference learning can be beneficial. Instead of predicting $x_{t+1}$ directly, we predict the increment $\Delta x_{t}=x_{t+1}-x_{t}$. This approach, often encountered in reinforcement learning (sutton2018reinforcement), can reduce biases from unobserved global offsets or stable large-scale means. For instance, daily vegetation changes or seasonal fluctuations in species population can be smaller and more stable than absolute population numbers. By focusing on differences, we encourage the model to learn transition dynamics or, more specifically, biodiversity dynamics:
[…]
This choice is reinforced by empirical results in specific biodiversity modelling tasks (e.g., ephemeral wetlands or short-lived insect populations), which show improved forecasting stability over standard next-step MSE (clark2001ecological).
@@ 3.4 Fine-tuning
BioAnalyst is fine-tuned in two different settings, each contributing to a distinct goal. In this section, a description of each setting is provided, and in the next section, we present the quantitative results.
@@
Roll-out finetuning: In this setting, we fine-tune the entire BioAnalyst for six and twelve rollout steps, effectively predicting biodiversity dynamics six months and one year ahead. Fine-tuning with a mix of variable horizons is a technique successfully used on Aurora (bodnar2024aurora), PrithviWxC (schmude2024prithviwxcfoundationmodel) and Aardvak (allen2025end), enhancing long-horizon forecasting capabilities. In practice, we freeze the whole architecture, including the encoder, decoder, and backbone, while training only the newly added VeRA adapters (kopiczko2024vera) on the backbone’s attention heads. We found VeRA to perform equally or sometimes slightly better than other Parameter Efficient Fine-tuning Techniques (PEFTs), such as LoRA (hu2022lora), which use only one-tenth of the learnable parameters. More specific, starting from an observed state $x_{t}$, the model is unrolled autoregressively: at each step $k$ it takes the previous prediction $\hat{x}_{t+k-1}$ as input and produces a new prediction $\hat{x}_{t+k}$. We supervise all intermediate steps with the same task-dependent loss $\mathcal{L}_{TD}$ and define the $K$-step rollout loss as
$\mathcal{L}_{\text{rollout}}(t,K)=\frac{1}{K}\sum_{k=1}^{K}\mathcal{L}_{TD}\!\left(\hat{x}^{u}_{t+k},\,x^{u}_{t+k}\right).$ (10)
Our mix horizons are $K=6$ and $K=12$ months. To keep training stable and affordable for long sequences, we follow the “pushforward trick” of brandstetter2022message: during backpropagation we stop gradients at the inputs of steps $k>1$, so that the loss is averaged over all steps for evaluation, but only the final step ($k=K$) contributes gradients to the model parameters.
@@
Task specific finetuning: To evaluate the ecological capacity and environmental structure encoded in BioAnalyst, we implement two complementary fine-tuning tasks. These tasks are designed to interrogate distinct dimensions of the model’s representation space: its ability to adapt to biotic presence-only data under temporal shift, and its capacity to retain structured abiotic gradients related to seasonal climate. For comparison, we also benchmark the performance on these fine-tuning tasks using four baselines and the Aurora-0.25 model (bodnar2024aurora), which shares the same model architecture but has been pre-trained on climate modalities only. More specific, the first is a joint species distribution task that involves partial model adaptation: the BioAnalyst’s encoder and decoder are frozen. At the same time, the backbone is fine-tuned with VeRA adapters using historical (2017-2021) species vascular plant occurrence (500 most commonly occurring species) data from the GeoLifeCLEF2024 benchmark dataset (geolifeclef-2024) to forecast time-series distributions. The second task is diagnostic: a regression head is trained on top of frozen decoder embeddings for both BioAnalyst and Aurora, to predict monthly climate variables (2-meter temperature and precipitation) from CHELSA v2.1 (karger2017climatologies) (2000-2019). Together, these tasks provide a dual lens on the model’s ecological generalisation and environmental coherence. The detailed information for the model card, software, training and scaling, finetuning tasks and metrics is available at Appendix C.
@@ 4.1 Rollouts: Forecasting biodiversity dynamics
Figure 3: Mean Absolute Error for the 28 animal species on a 12-step rollout. Blue line highlights the performance of roll-out finetuned BioAnalyst for $K=6$ steps while orange line for $K=12$ steps.
In this part, we present four main results produced from our pre-training and rollout fine-tuning experiments whose target is to model and forecast biodiversity dynamics. First we present the MAE for the species variable group (28 animal species), comparing the performance of a 12-step roll-out for our mix horizons $K=6$ and $K=12$ time-steps respectively Figure 3. The 12-step finetuned variant follows a lower MAE across the 12-step trajectory, validating the effectiveness of our mix horizon scheme during roll-out finetuning. Both variants exhibit increasing MAE after 10 steps. Second, we report a mean Sørensen similarity of 0.31 for 28 animal species across all evaluated land grid cells, indicating that the predicted assemblages recover roughly one-third of the observed community composition. Figure 4 highlights spatial variation in assemblage agreement, with higher similarity in western and central Europe and lower values in eastern and south-eastern regions. This pattern closely follows known spatial biases in the underlying GBIF occurrence data, with denser sampling in some regions than others, so the map reflects both model skill and uneven observation effort. Overall, the spatial pattern suggests moderate compositional skill in well-sampled areas, useful for broad biogeographic inference, yet leaving room for improving species co-occurrence dynamics. Third, Figure 5 depicts BioAnalyst ability to localise and spatially predict granular species distributions. Fourth, our investigation on BioAnalysts cross-attention weights, highlighted that climate variables receive the highest attention, followed by the species and surface variables, while most other modality groups receive lower attention. Spatial analysis, particularly in the case of a high-interest variable (i.e., species), reveals that attention is not uniformly distributed, rather than concentrates on Scandinavia and Central Europe regions Appendix D.
Figure 4: Community Sørensen similarity between predicted and observed plant assemblages for 28 animal species across Europe, based on GBIF presence records. The mean similarity is $\bar{S}=0.31$, indicating that the model recovers roughly one third of the recorded community composition. Warm colours (yellow–red) denote higher assemblage agreement, while cool blues indicate little overlap and/or sparsely sampled regions.
Figure 5: (a) Ground truth and prediction spatial plots for the species Pieris brassicae (ID 1920506) on 01-4-2019 and MAE of 0.00003. (b) Zoomed in plot, highlighting areas of interest, where the model can capture the general distribution, although unable to capture high-density areas.
@@ 4.2 Biotic fine-tuning: forecasting species distributions
To assess the biotic predictive capacity of BioAnalyst’s pre-trained embeddings, we fine-tuned a classification head to forecast species presence-absence for 500 vascular plant species in the GeoLifeCLEF 2021 dataset. The model achieved high accuracy with an F1 score of 0.9964 and an RMSE of 0.5284, surpassing our selected baselines LatentMLP and ConvLST, demonstrating its ability to support macro-scale joint species distribution modelling. In the same setting, Aurora performs marginally worse in evaluation metrics than BioAnalyst, as shown in Table 1. Nevertheless, investigating the predicted species richness in more detail for both models, we observe that Aurora produces a more spatially extensive richness field, assigning moderate richness also to cells with recorded absent observations, which likely reflects extrapolation from smooth climate fields. BioAnalyst’s predictions remain more tightly constrained to the observed footprint, consistent with the additional process-related modalities it uses (land use, vegetation, species signals) Figure 6. This result supports the view that, for complex community-level processes, climate-only representations can be limiting, and that integrating additional process-related modalities, as in BioAnalyst, can substantially improve predictive performance and spatial realism.
Figure 6: Comparison of observed and predicted species richness in the Netherlands for the year 2021. a) The Observed Richness based on empirical field data from GeoLifeCLEF2024 presence-absence survey; blank cells represent true zeros indicating absence of species occurrence. b) The species richness for 2021 predicted by the BioAnalyst, and c) Aurora-0.25 predictions for 2021.
Table 1: Performance of the species distribution forecasting and of the linear probing task. Note: $R^{2}$ is not applicable to this classification-style species forecasting task and therefore not reported.
Model species distribution forecast linear probing
Loss F1 RMSE Loss $R^{2}$ RMSE
BioAnalyst 0.0057 0.9964 0.5284 0.0225 0.9002 0.1499
Aurora 0.0130 0.9945 0.5014 0.2668 0.7354 0.5144
RandomForest 0.7260 0.2426
SVM 0.1741 0.3828
LatentMLP 0.0435 0.8916 0.6951
ConvLSTM 0.0256 0.9924 0.5433
@@ 4.3 Abiotic linear probing: recovering seasonal climate structure
Figure 7: Comparison between CHELSA target climate fields and decoded outputs from the BioAnalyst decoder for September 2011 over Europe. Panels (a) and (b) show the downsampled CHELSA reference for precipitation (kg $m^{2}$) and temperature (K/10), respectively. Panels (c) and (d) show the corresponding decoded predictions from the model after reconstruction.
The model achieved strong predictive performance, with the best epoch reaching an $R^{2}$ of 0.9002, a loss of 0.0225, and an RMSE of 0.1499, as shown in Table 1. These values indicate that the BioAnalyst decoder outputs contain sufficient information to reconstruct fine-grained seasonal climate patterns even without any fine-tuning of the encoder Figure 7. The decoded precipitation captures major spatial gradients and orographic patterns (e.g., Alps, Norway), though with some smoothing. The decoded temperature field accurately reproduces latitudinal and coastal gradients present in the target. In the same theme, comparing BioAnalyst with baselines such as the pre-trained Aurora-025, a Random Forest model and Support Vector Machine yielded a stronger predictive capacity for BioAnalyst. However, this comparison should be interpreted only as an indicative baseline: BioAnalyst is pre-trained directly on monthly biodiversity states, whereas Aurora is an Earth-system foundation model that is pre-trained to forecast geophysical variables at short lead times (typically 6-hour steps) and fine-tuned for weather, wave and air-quality tasks rather than biodiversity Appendix C.4.4.
@@ 5 Conclusion & Discussion
In this work, we introduced BioAnalyst, to our knowledge the first Foundation Model for biodiversity. BioAnalyst is truly multi-modal, light-weight and can be used to model various complex spatio-temporal ecological phenomena with competitive performance for regional and nation-scale applications at $0.25^{\circ}$ resolution, setting a new accuracy benchmark for ecological forecasting. We highlighted its predictive analytics as a stand-alone model in tasks such as biodiversity dynamics modelling, as well as in tasks like joint species distribution forecasting, absence detection, and monthly climate linear probing. This opens new opportunities for rapid model adaptation in data-poor contexts and for advancing hypothesis-driven ecological modelling through representation learning.
However, some caution is warranted when interpreting downstream predictions. For instance, while BioAnalyst captures species occurrence and absence patterns with adequate accuracy (Figure 6), this success may partly reflect temporal biases in observation effort rather than actual ecological change. Although this modelling setup focuses on presence–absence, the multi-taxon nature of the dataset effectively means we are performing joint species distribution modelling (jSDM) across both space and time, which is an notoriously difficult task in classical ecology tikhonov2020joint; konig2021scale. Also, BioAnalyst, comes with a series of limitations like the lack of uncertainty quantification of the predictions, the constrained area of operation, Europe, which is not representative of global biodiversity dynamics. Since BioAnalyst is trained on data that represent biodiversity only in the terrestrial realm, it does not take into consideration the seas/oceans/lakes/dykes, for example, amplifying the bias towards other parts of biodiversity and the compound effects they may have on both regional and national biodiversity dynamics. Applications like local habitat management and single-site reserve design will require downscaling or complementary, higher-resolution data for further fine-tuning. Still, there are many open challenges ahead, one of the most difficult to tackle is data availability and methods to improve ecology-based downstream tasks.
@@
Part of the future work will include more frequently sampled, higher resolution and better-curated data points that can further enhance BioAnalyst’s capabilities in combination with longer training. A user interaction interface could enhance BioAnalyst’s interpretability and ease of use. Methods for quantifying this uncertainty at different granularity levels, such as cartograms (ROCCHINI2019146) or meta-model traits (okanik2024uncertainty), are planned as BioAnalyst’s future work. Truly synergistic models that embed ecological principles (e.g. energy budgets, trophic interactions) into AI architectures that simulate population dynamics are an exciting frontier (agarwal2025generalgeospatialinferencepopulation). This could mean neural networks that respect mass-balance constraints or reinforcement learning agents that simulate animal foraging behaviour. Such integration would yield models that not only predict well but also adhere to known ecological laws, making them more generalisable, trustworthy, and aligned with global biodiversity goals (DESANTIS2025100198).
Reproducibility statement
All our code is open-sourced under the MIT license and can be found at:
Data availability: The dataset used can be found at https://huggingface.co/datasets/BioDT/BioCube and its code base at https://github.com/BioDT/bfm-data.
For the complete pre-training and task fine-tuning data information, please look at Appendix B. The model card, pre-training and fine-tuning configurations are available at Appendix C. Additional ablation studies and interpretability material are available at Appendix D
Author Contributions
Acknowledgments